# Hyperparameter Transfer with Mixture-of-Experts Layers

### v2

@@ D.2 Expert load balancing
for all parameters groups, so that cross-validation loss and load balancing regularization are on the same scale. However, the practical implications may be more complex as (a) the router matrix at each layer is predominantly responsible for the load balancing at this layer, and it is not clear whether it is desirable or not for FFN or self-attention modules to receive the same gradient norm from load balancing (in other layers) as they would from cross-entropy, or if one router matrix in some layer needs to receive gradients from load-balancing other layers. One possible remedy is to restrict per-layer load balancing gradients to only the router matrix of that specific layer, but this seems to deviate from known practices nonetheless.
Furthermore, under the standard Switch-transformer (Fedus et al., 2022) framework with softmax activation, it’s possible for a phase transition where router logits (expert mixing coefficients) go from near uniform (at initialization) to close to one-hot, making designing the desiderata for HP transfer and $\alpha$ scale tricky. Whereas in our case, the un-normalized mixing coefficients $g$’s through a sigmoid gate are always bounded and well-behaved.
Finally, when we finally justify HP transfer on training or validation loss, usually the object of study is only concerning cross-entropy and not with the regularized load-balancing loss. It has further been reported (Wang et al., 2024) in some cases that introducing too much auxiliary loss harms performance on cross-entropy loss. The tradeoff for quicker load balance versus validation loss deviates from the pure study of HP transfer and we do not pursue here.
@@ D.3 More experiments
Figure 17: Maximum load imbalance in our class of scaling models. Load imbalance (Eqn. 2) is calculated as the max, over $L\cdot n_{\operatorname{embd}}$ total experts in the model, absolute value of (proportion of tokens routed to the expert) subtracts (the target uniform sparsity $\kappa$), similar to Figure˜9. Here $y$-axis is from 0 to $1-\kappa$. We observe that load balancing convergence is perfect when we increase width and expert size, the number of experts per layer, but may be compromised when we scale up depth for too large, which we hypothesize is due to constant-HP tuning issues and may require further investigation (the same hypothesis bars us from reproducing reliable loss collapse in early steps for Figure˜3 when scaling number of layers). The last column shows that load balance convergence persists even when training horizon is extended.
@@ Appendix E Dynamical mean field theory
resulting in the following $\mathcal{O}(1)$ action $\mathcal{S}$ where we suppress data and time indices
@@
$+\nu\sum_{\ell}\left[\widehat{M}^{\ell}_{\sigma\sigma C_{\phi}}{M}^{\ell}_{\sigma\sigma C_{\phi}}+\widehat{M}^{\ell}_{\sigma\sigma C_{g}}{M}^{\ell}_{\sigma\sigma C_{g}}+\widehat{M}^{\ell}_{\dot{\sigma}\dot{\sigma}\mathcal{A}\mathcal{A}}{M}^{\ell}_{\dot{\sigma}\dot{\sigma}\mathcal{A}\mathcal{A}}-\widehat{\bar{R}}^{\ell}_{\phi\xi}\bar{R}^{\ell}_{\phi\xi}-\widehat{\bar{R}}^{\ell}_{g\chi}\bar{R}^{\ell}_{g\chi}\right]$
$+\ln\mathcal{Z}_{\text{res}}+\nu\sum_{\ell=1}^{L}\ln\mathcal{Z}_{\text{exp}}^{\ell}$ (27)
@@
$\exp\left(-\frac{i}{N_{e}}\widehat{\bar{R}}_{\phi\xi}^{\ell}\sigma_{k}^{\ell}\phi(h^{\ell}_{k})\widehat{\xi}^{\ell}_{k}-\frac{i}{N_{e}}\widehat{\bar{R}}_{g\chi}^{\ell}\sigma_{k}^{\ell}g^{\ell}_{k}\widehat{\chi}^{\ell}_{k}\right)$ (31)
We let $N_{e}(N)$ and $E(N)$ be diverging functions for the hidden width and expert size as a function of residual stream width $N$ at any fixed value of depth $L$. We assume the following condition to be satisfied
$\lim_{N\to\infty}\frac{N}{N_{e}(N)E(N)}=\alpha_{\star}L<\infty$ (32)
@@
Large Expert Size
For any diverging expert size $N_{e}\to\infty$ we can expand $\mathcal{Z}^{\ell}_{\text{within-exp}}$ to obtain a reduced action
$N_{e}\ln\mathcal{Z}_{\text{within-exp}}^{\ell}$ $\sim-\widehat{C}^{\ell}_{\phi_{k}}\{\phi(h_{k}^{\ell})\phi(h^{\ell}_{k})\}-\widehat{C}^{\ell}_{g_{k}}\{g^{\ell}_{k}g^{\ell}_{k}\}-\widehat{C}^{\ell}_{\phi\xi}\{\phi(h_{k}^{\ell})\xi^{\ell}_{k}\}$
$-i\widehat{\bar{R}}^{\ell}_{\phi\xi}\sigma_{k}^{\ell}\{\phi(h^{\ell}_{k})\widehat{\xi}^{\ell}_{k}\}-i\widehat{\bar{R}}^{\ell}_{g\chi}\sigma_{k}^{\ell}\{g^{\ell}_{k}\widehat{\chi}^{\ell}_{k}\}+\mathcal{O}(N_{e}^{-1})$ (33)
where $\{\}$ represents an average over the following stochastic process defined by (which is a function of all of the per-expert order parameters) and has the structure
$h^{\ell}_{k}=\chi^{\ell}_{k}+\gamma_{0}\Delta\sigma_{k}^{\ell}C_{h}^{\ell}g^{\ell}_{k}\ ,\ \chi^{\ell}_{k}\sim\mathcal{GP}(0,C_{h}^{\ell})$
$g^{\ell}_{k}=\dot{\phi}(h^{\ell}_{k})\left[\xi^{\ell}_{k}+\gamma_{0}\Delta\sigma_{k}^{\ell}C_{g}^{\ell+1}\phi(h^{\ell}_{k})\right]\ ,\ \xi^{\ell}_{k}\sim\mathcal{GP}(0,C_{g}^{\ell+1})$ (34)
These equations provide the within-neuron distribution and dynamics as $N_{e}\to\infty$. We note that the response functions from the residual stream have no influence on these dynamics in this limit.
Large Residual Stream + Large Number of Experts
@@
$\frac{\partial\mathcal{S}}{\partial\widehat{M}_{\sigma\sigma C_{\phi_{k}}}^{\ell}}$ $=M_{\sigma\sigma C_{\phi_{k}}}^{\ell}-\left[\sigma^{\ell}\sigma^{\ell}\{\phi(h_{k}^{\ell})\phi(h_{k}^{\ell})\}\right]=0$ (37)
$\frac{\partial\mathcal{S}}{\partial\widehat{M}_{\sigma\sigma C_{g_{k}}}^{\ell}}$ $=M_{\sigma\sigma C_{g_{k}}}^{\ell}-\left[\sigma^{\ell}\sigma^{\ell}\{g_{k}^{\ell}g_{k}^{\ell}\}\right]=0$ (38)
$\frac{\partial\mathcal{S}}{\partial\widehat{M}_{\dot{\sigma}\dot{\sigma}\mathcal{A}\mathcal{A}}^{\ell}}$ $=M_{\dot{\sigma}\dot{\sigma}\mathcal{A}\mathcal{A}}^{\ell}-\left[\dot{\sigma}^{\ell}\dot{\sigma}^{\ell}\mathcal{A}^{\ell}\mathcal{A}^{\ell}\right]=0$ (39)
$\frac{\partial\mathcal{S}}{\partial\widehat{\bar{R}}_{\phi\xi}^{\ell}}$ $=-\bar{R}^{\ell}_{\phi\xi}-i\left[\sigma^{\ell}\left\{\phi(h^{\ell}_{k})\widehat{\xi}^{\ell}_{k}\right\}\right]=0$ (40)
$\frac{\partial\mathcal{S}}{\partial\widehat{\bar{R}}_{g\chi}^{\ell}}$ $=-\bar{R}^{\ell}_{g\chi}-i\left[\sigma^{\ell}\left\{g^{\ell}_{k}\widehat{\chi}^{\ell}_{k}\right\}\right]=0$ (41)
$\frac{\partial\mathcal{S}}{\partial{\bar{R}}_{\phi\xi}^{\ell}}$ $=-\widehat{\bar{R}}^{\ell}_{\phi\xi}-i\left<\widehat{\bar{\chi}}^{\ell}g^{\ell}\right>=0$ (42)
$\frac{\partial\mathcal{S}}{\partial{\bar{R}}_{g\chi}^{\ell}}$ $=-\widehat{\bar{R}}^{\ell}_{g\chi}-i\left<\widehat{\bar{\xi}}^{\ell}h^{\ell}\right>=0$ (43)
@@
Final DMFT Equations
The DMFT single site equations are the following. For the residual stream (reminder that $\alpha_{\star}=\lim_{N\to\infty}\frac{N}{N_{e}(N)E(N)L}$ is bounded).
@@
$h^{\ell}_{k}(\bm{x},t)=\chi^{\ell}_{k}(\bm{x},t)+\gamma_{0}\mathbb{E}_{\bm{x}^{\prime}}\int dt^{\prime}\Delta(\bm{x}^{\prime},t^{\prime})\sigma_{k}^{\ell}(\bm{x}^{\prime},t^{\prime})C_{h}^{\ell}(\bm{x},\bm{x}^{\prime},t,t^{\prime})g_{k}(\bm{x}^{\prime},t^{\prime})\ ,\ \chi^{\ell}_{k}(\bm{x},t)\sim\mathcal{GP}(0,C_{h}^{\ell})$
$z^{\ell}_{k}(\bm{x},t)=\xi^{\ell}_{k}(\bm{x},t)+\gamma_{0}\mathbb{E}_{\bm{x}^{\prime}}\int dt^{\prime}\Delta(\bm{x}^{\prime},t^{\prime})\sigma_{k}^{\ell}(\bm{x}^{\prime},t^{\prime})C_{g}^{\ell+1}(\bm{x},\bm{x}^{\prime},t,t^{\prime})\phi(h_{k}(\bm{x}^{\prime},t^{\prime}))\ ,\ \xi^{\ell}_{k}(\bm{x},t)\sim\mathcal{GP}(0,C_{g}^{\ell+1})$
@@
The average $\{\}$ over neurons within each expert are determined by the above stochastic process.
@@
For $\alpha_{\star}=0$, the Brownian motion term disappears and the residual stream dynamics reduce to a neural ODE, consistent with CompleteP scaling (Dey et al., 2025; Chizat, 2025).
E.1 Verifying Convergence in Soft Routing Model

### v3

@@ D.2 Expert load balancing
for all parameters groups, so that cross-validation loss and load balancing regularization are on the same scale. However, the practical implications may be more complex as the router matrix at each layer is predominantly responsible for the load balancing at this layer, and it is not clear whether it is desirable or not for FFN or self-attention modules to receive the same gradient norm from load balancing (in other layers) as they would from cross-entropy, or if one router matrix in some layer needs to receive gradients from load-balancing other layers. For instance, should the MLP in Layer 1 be responsible (in terms of gradient scale) for load balancing in Layer 10? Furthermore, when we finally justify HP transfer on training or validation loss, usually the object of study is only concerning cross-entropy and not with the regularized load-balancing loss. So the tradeoff between better validation loss versus strict uniform balance in early training (Dai et al., 2024; Liu et al., 2024) cannot be evaluated fairly without making restricting assumptions.
In conclusion, analyzing a global auxiliary loss requires much more practitioner-specific assumptions (e.g., desiderata in balancing gradients in different layers, desiderata in comparing different models), and that blanket application is not theoretically justified (in fact, aux-loss is not globally minimized with balanced routing). Furthermore, even the empirical application of aux-loss requires scrutiny (e.g. batch-wise vs sequence-wise balancing achieve different objectives (Liu et al., 2024)). While still not a perfect solution, our choice of aux-loss-free balancing is both empirically popular as a replacement and rids us of having to make much more fine-grained assumptions (e.g. only specific biases receive balancing gradients, and sequence vs batch imbalance losses coincide).
@@ D.3 More experiments
  8. 8.
    In Figure˜18, we test and verify that softmax vs sigmoid activation does not significantly (if at all) impact our qualitative transfer observations.
@@
Figure 17: Maximum load imbalance in our class of scaling models. Load imbalance (Eqn. 2) is calculated as the max, over $L\cdot n_{\operatorname{embd}}$ total experts in the model, absolute value of (proportion of tokens routed to the expert) subtracts (the target uniform sparsity $\kappa$), similar to Figure˜9. Here $y$-axis is from 0 to $1-\kappa$. The last column shows that load balance convergence persists even when training horizon is extended.
Figure 18: Transfer behavior on Fineweb when sigmoid expert mixing is replaced with softmax. Our analysis remains applicable and empirically verified.
@@ Appendix E Dynamical mean field theory
Universal Dynamics in A Joint Scaling Limit
We will start our analysis by assuming that all variables $N,E,N_{e},L$ diverge together with (1) constant sparsity $K=\kappa E$ and (2) the following condition
$\lim_{N\to\infty}\frac{N}{N_{e}(N)L(N)E(N)}\equiv\alpha_{\star}<\infty$ (5)
We will first establish that the DMFT equations in terms of arbitrary finite $\alpha_{\star}$. However, since most commonly utilized joint scaling protocols result in $\alpha_{\star}=0$ we will then specialize to that case. In that case, we show that the limit dynamics follow a universal neural ODE that depends on $\kappa$ but not on the FFN ratio $N_{e}/N$.
@@
resulting in the following $\mathcal{O}(1)$ action $\mathcal{S}$ where we suppress data and time indices, where we let $\nu=E/N$ and $\alpha_{\star}=\frac{N}{N_{e}EL}$
@@
$+\nu\sum_{\ell}\left[\widehat{M}^{\ell}_{\sigma\sigma C_{\phi}}{M}^{\ell}_{\sigma\sigma C_{\phi}}+\widehat{M}^{\ell}_{\sigma\sigma C_{g}}{M}^{\ell}_{\sigma\sigma C_{g}}+\widehat{M}^{\ell}_{\dot{\sigma}\dot{\sigma}\mathcal{A}\mathcal{A}}{M}^{\ell}_{\dot{\sigma}\dot{\sigma}\mathcal{A}\mathcal{A}}\right]-\frac{1}{\alpha_{\star}L}\sum_{\ell}\left[N_{e}\widehat{\bar{R}}^{\ell}_{\phi\xi}\bar{R}^{\ell}_{\phi\xi}+N_{e}\widehat{\bar{R}}^{\ell}_{g\chi}\bar{R}^{\ell}_{g\chi}\right]$
$+\ln\mathcal{Z}_{\text{res}}+\nu\sum_{\ell=1}^{L}\ln\mathcal{Z}_{\text{exp}}^{\ell}$
$\nu=\frac{E}{N}\ ,\ \alpha_{\star}=\frac{N}{N_{e}EL}$ (28)
@@
$\exp\left(-i\widehat{\bar{R}}_{\phi\xi}^{\ell}\sigma_{k}^{\ell}\phi(h^{\ell}_{k})\widehat{\xi}^{\ell}_{k}-i\widehat{\bar{R}}_{g\chi}^{\ell}\sigma_{k}^{\ell}g^{\ell}_{k}\widehat{\chi}^{\ell}_{k}\right)$ (32)
We let $N_{e}(N)$ and $E(N)$ be diverging functions for the hidden width and expert size as a function of residual stream width $N$ at any fixed value of depth $L(N)$ (possibly constant or diverging). We assume the following condition to be satisfied
$\lim_{N\to\infty}\frac{N}{N_{e}(N)E(N)L(N)}=\alpha_{\star}=0$ (33)
@@
Saddle Point Equations
@@
$\frac{\partial\mathcal{S}}{\partial\widehat{M}_{\sigma\sigma C_{\phi_{k}}}^{\ell}}$ $=\nu M_{\sigma\sigma C_{\phi_{k}}}^{\ell}-\nu\left[\sigma^{\ell}\sigma^{\ell}\{\phi(h_{k}^{\ell})\phi(h_{k}^{\ell})\}\right]=0$ (36)
$\frac{\partial\mathcal{S}}{\partial\widehat{M}_{\sigma\sigma C_{g_{k}}}^{\ell}}$ $=\nu M_{\sigma\sigma C_{g_{k}}}^{\ell}-\nu\left[\sigma^{\ell}\sigma^{\ell}\{g_{k}^{\ell}g_{k}^{\ell}\}\right]=0$ (37)
$\frac{\partial\mathcal{S}}{\partial\widehat{M}_{\dot{\sigma}\dot{\sigma}\mathcal{A}\mathcal{A}}^{\ell}}$ $=\nu M_{\dot{\sigma}\dot{\sigma}\mathcal{A}\mathcal{A}}^{\ell}-\nu\left[\dot{\sigma}^{\ell}\dot{\sigma}^{\ell}\mathcal{A}^{\ell}\mathcal{A}^{\ell}\right]=0$ (38)
$\frac{\partial\mathcal{S}}{\partial\widehat{\bar{R}}_{\phi\xi}^{\ell}}$ $=-\nu\bar{R}^{\ell}_{\phi\xi}-i\nu\left[\sigma^{\ell}\left\{\phi(h^{\ell}_{k})\widehat{\xi}^{\ell}_{k}\right\}\right]=0$ (39)
$\frac{\partial\mathcal{S}}{\partial\widehat{\bar{R}}_{g\chi}^{\ell}}$ $=-\nu\bar{R}^{\ell}_{g\chi}-i\nu\left[\sigma^{\ell}\left\{g^{\ell}_{k}\widehat{\chi}^{\ell}_{k}\right\}\right]=0$ (40)
$\frac{\partial\mathcal{S}}{\partial{\bar{R}}_{\phi\xi}^{\ell}}$ $=-\frac{1}{\alpha_{\star}L}\widehat{\bar{R}}^{\ell}_{\phi\xi}-i\left<\widehat{\bar{\chi}}^{\ell}g^{\ell}\right>=0$ (41)
$\frac{\partial\mathcal{S}}{\partial{\bar{R}}_{g\chi}^{\ell}}$ $=-\frac{1}{\alpha_{\star}L}\widehat{\bar{R}}^{\ell}_{g\chi}-i\left<\widehat{\bar{\xi}}^{\ell}h^{\ell}\right>=0$ (42)
@@
We note that the functions $\widehat{\bar{R}}^{\ell}_{\phi\xi}$ and $\widehat{\bar{R}}^{\ell}_{g\chi}$ are both $\mathcal{O}\left(\frac{N}{N_{e}EL}\right)=\mathcal{O}(\alpha_{\star})$ [^3] To see this compute $\frac{\partial h}{\partial r}$ or $\frac{\partial g}{\partial u}$ and see that they are $\sim\mathcal{O}(L^{-1})$. Using the fact that $\frac{1}{\nu N_{e}}=\frac{N}{N_{e}E}$, this verifies the correct scale of $\widehat{\bar{R}}$. We thus introduce the following rescaled definitions to simplify our expression
$\widehat{\bar{R}}^{\ell}_{\phi\xi}\equiv\alpha_{\star}A^{\ell}\ ,\ \widehat{\bar{R}}^{\ell}_{g\chi}\equiv\alpha_{\star}B^{\ell}$ (44)
We will express the final single site equations in terms of $A^{\ell}$ and $B^{\ell}$ which are dimensionless at leading order.
@@
DMFT Equations
The DMFT single site equations under the assumption that
$\alpha_{\star}\equiv\frac{N}{N_{e}LE},$ (47)
are the following for the the residual stream
@@
$h^{\ell}_{k}(\bm{x},t)=u^{\ell}_{k}(\bm{x},t)+\alpha_{\star}\int dtd\bm{x}^{\prime}A^{\ell}(\bm{x},\bm{x}^{\prime},t,t^{\prime})g^{\ell}_{k}(\bm{x}^{\prime},t^{\prime})$ (50)
$+\gamma_{0}\mathbb{E}_{\bm{x}^{\prime}}\int dt^{\prime}\Delta(\bm{x}^{\prime},t^{\prime})\sigma_{k}^{\ell}(\bm{x}^{\prime},t^{\prime})C_{h}^{\ell}(\bm{x},\bm{x}^{\prime},t,t^{\prime})g_{k}^{\ell}(\bm{x}^{\prime},t^{\prime})\ ,\ u^{\ell}_{k}(\bm{x},t)\sim\mathcal{GP}(0,C_{h}^{\ell})$
$z^{\ell}_{k}(\bm{x},t)=r^{\ell}_{k}(\bm{x},t)+\alpha_{\star}\int dtd\bm{x}^{\prime}B^{\ell+1}(\bm{x},\bm{x}^{\prime},t,t^{\prime})\phi(h^{\ell}_{k}(\bm{x}^{\prime},t^{\prime}))$ (51)
$+\gamma_{0}\mathbb{E}_{\bm{x}^{\prime}}\int dt^{\prime}\Delta(\bm{x}^{\prime},t^{\prime})\sigma_{k}^{\ell}(\bm{x}^{\prime},t^{\prime})C_{g}^{\ell+1}(\bm{x},\bm{x}^{\prime},t,t^{\prime})\phi(h_{k}^{\ell}(\bm{x}^{\prime},t^{\prime}))\ ,\ \xi^{\ell}_{k}(\bm{x},t)\sim\mathcal{GP}(0,C_{g}^{\ell+1})$
@@
The average $\{\}$ over neurons within each expert are determined by the above stochastic process. We see that the response terms involving $A^{\ell}$ and $B^{\ell}$
@@
For $\alpha_{\star}=0$, the Brownian motion term disappears and the residual stream dynamics reduce to a neural ODE, consistent with CompleteP scaling (Dey et al., 2025; Chizat, 2025). Further, for $\alpha_{\star}=0$ the response function terms for the within-FFN block neuron dynamics disappear, resulting in simpler within block dynamics. This analysis points to finite $\alpha_{\star}$ as worthy of future study as they provide an example of complete feature learning in a Neural SDE, which was previously thought to require an ODE limit (Bordelon et al., 2024; Dey et al., 2025).
Key Finding: FFN Ratio Indepedence
The ratio of the FFN width $N_{e}/N$ only appears in the ratio $\alpha_{\star}$. Provided this variable converges rapidly to zero as the model approaches the limit, the limit dynamics will not depend on the FFN ratio. In practice, most of the common joint scaling strategies scale depth and expert width linearly with the residual width, which results in $\alpha_{\star}\sim\frac{1}{N}\to 0$.
E.1 Error Analysis
In this section, we discuss the asymptotic errors to the joint scaling limit. We are interested in quantifying the scale of the error in network outputs $f$ at finite $N,E,N_{e},L$ and the joint infinite limit with $\alpha_{\star}=\frac{N}{EN_{e}L}\to 0$. Concretely, we aim to bound
$|f(N,E,N_{e},L)-f_{\infty}|.$ (56)
The key sources of finite size errors near the asymptotic limit are
  1. 1.
    Finite Residual Stream Fluctuations: Imperfect concentration of residual stream kernels $C_{h}$ and $C_{g}$ induce central limit theorem fluctuations on scale $\mathcal{O}(N^{-1/2})$. These fluctuations also impact the model outputs.
  2. 2.
    Finite Expert Fluctuations Imperfect concentration of expert averaged kernels $M_{\sigma\sigma C_{\phi}},M_{\sigma\sigma C_{g}}$, etc. induce errors on the order of $\mathcal{O}(E^{-1/2})$.
  3. 3.
    Finite Expert Size Corrections The finite expert size effects contribute in two places. First, they impact the residual stream dynamics at leading order at a scale $\mathcal{O}(E^{-1/2}N_{e}^{-1/2}+N_{e}^{-1})$ as they perturb $M$ kernels. However, since these kernels are also expert averages, the CLT fluctuations scale as for large $E$ (mean-zero fluctuations from finite $N_{e}$ in the expert averages $\{\}$ average out to zero when we compute averages over the expert population $\left[\right]$). Second, the within-expert response functions at finite $\alpha_{\star}$ perturb the within-expert dynamics by a scale $\alpha_{\star}=\frac{N}{N_{e}EL}$.
  4. 4.
    Noise on Residual Stream vs Neural ODE The contribution from the $u^{\ell}$ and $r^{\ell}$ variables contribute variance on the scale $\alpha_{\star}=\frac{N}{LN_{e}E}$. These cause an error in the predictor at a scale of $\sim\alpha_{\star}$. We note that our error metric is focused on the error in the kernels and model outputs, not on the residual stream itself which would have $\sqrt{\alpha_{\star}}$ as the error.
  5. 5.
    Finite Depth Discretization The kernels at large depth $L$ can be viewed as an Euler discretization $C_{h}^{\ell+1}=C_{h}^{\ell}+\frac{1}{L}[$...] of an underlying ODE. This discretization induces an error of order $\sim 1/L$.
In total we arrive at an error bound of the form
$|f(N,E,N_{e},L)-f_{\infty}|=\mathcal{O}\left(\frac{1}{\sqrt{N}}+\frac{1}{\sqrt{E}}+\frac{1}{N_{e}}+\frac{N}{N_{e}EL}+\frac{1}{L}\right).$ (57)
The approximation error is minimized by taking the following joint scaling as the residual stream width $N$ diverges
$E\sim N\ ,\ L\sim\sqrt{N}\ ,\ N_{e}\sim\sqrt{N}$ (58)
Under this regime $\alpha_{\star}\sim\frac{1}{N}\to 0$ as desired. If we define the total parameters $P=NEN_{e}L\sim N^{3}$ then the scaling law for the error should be $\mathcal{O}(P^{-1/6})$.
E.2 Verifying Convergence in Soft Routing Model